# BioMoDiffuse: Physics-Guided Biomechanical Diffusion for Controllable and Authentic Human Motion Synthesis

### v1

@@ 3.5 Pose Refinement with Diffusion Inversion
Figure 3: Refining pose using our method. Building on existing pose estimation models, we integrate BioMoDiffuse with DDIM inversion [62]. By adopting an inversion-then-forward pipeline, our approach significantly improves the smoothness of generated SMPL sequences while enhancing the fitting accuracy of pose estimation.
@@
Our proposed framework can be readily adapted to refine pose estimates from existing pose estimation methods. The key insight is that our physics-guided diffusion model is capable of generating physically plausible motions starting from latent noise. Given the initial estimates of human poses, we first employ diffusion inversion—a method that reverses the usual diffusion process—to recover the latent noise corresponding to the poses. We then feed this noise into the physics-guided diffusion model, which injects biomechanical constraints to produce the refined poses that are more authentic compared to original estimates. Specifically, given the initial SMPL [39] parameters $\mathbf{x}_{0}$ from pose estimation, we first invert them to timestep $t$ using DDIM sampling:
@@
where $t$ controls the balance between information preservation and noise introduction. The inverted latent $\mathbf{x}_{t}$ then undergoes forward diffusion with our physics guidance to produce the refined poses:
@@
Crucially, our parametric implementation inherits the physics constraints from Secs. 3.2 and 3.3 without the need for additional tuning, demonstrating a unified biomechanical framework across both motion generation and refinement tasks.
4 Experiments
4.1 Datasets and Metrics
We conducted training and testing on our model using both the KIT-ML [51] and HumanML3D [14] datasets. The KIT-ML dataset contains 3,911 motions and 6,363 natural language annotations, while HumanML3D is derived from and reannotates the HumanAct12 [13] and AMASS [44] datasets, comprising a total of 14,616 motion sequences and 44,970 descriptions. We follow the evaluation metrics used in previous motion generation works [66, 82, 22], including Frechet Inception Distance (FID), Precision, Diversity, Multimodality, and Multi-Modal Distance. Additionally, we consider physics-related metrics such as skating ratio, ground floating, ground penetration, and foot clipping, as used in prior studies [18, 75, 77]. Furthermore, we introduce two novel metrics based on Fourier analysis: Velocity Consistency (VC) and Acceleration Consistency (AC). Our motivation stems from the observation that while the Smoothness metric captures global temporal coherence, it overlooks the inherent kinematic relationship between stable body parts and agile extremities in human motion. For a motion sequence represented as $\Phi\in\mathbb{R}^{T\times 3d}$, the metrics are formally defined as:
4.2 Quantitative Results
Figure 4: User Study Results.
User Study.
We asked participants to rate four aspects of the generated motions on a 1–5 scale: “Generation Quality,” “Motion Smoothness,” “Physical Plausibility,” and “Text Alignment.” Using the HumanML3D test set. Thirty users were invited to compare the motion generation results of our model, StableMoFusion, and Ground Truth across 20 prompts. The results demonstrate that our model exhibits superior performance in terms of generation quality, motion smoothness, physical plausibility, and semantic alignment compared to StableMoFusion. Furthermore, on certain examples, our method achieves results that are comparable to the ground truth, showcasing its competitive performance.
Table 2: Quantitative evaluation of physical plausible metrics on the HumanML3D and KIT-ML test set. All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables. The results for T2M, and MotionDiffuse on HumanML3D is sourced from ReinDiffuse [18].
@@ 4.2 Quantitative Results
Ours $\bm{0.135}$ $\bm{0.959}$ $0.385$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $0.002$
@@
Tab. 1 and Tab. 2 compare our method with existing approaches across both physical plausibility metrics (e.g., smoothness, foot sliding) and traditional generation metrics (FID, Diversity). Our model achieves state-of-the-art results in biomechanical authenticity while simultaneously advancing traditional motion quality, outperforming prior works across multiple dimensions. Experiments reveal critical limitations in existing paradigms: GPT-style generative models tend to over-smooth motions, failing to capture realistic dynamics, while diffusion-based methods exhibit temporal discontinuity—likely due to their implicit treatment of spatiotemporal relationships without explicit temporal modeling. By integrating our physics-aware network into the diffusion framework, we address these issues: the network’s frame-wise acceleration recalculation enhances temporal dependency learning, directly improving physical metrics. Concurrently, the introduction of biomechanical priors reduces solution space complexity, enabling superior performance in traditional metrics without compromising motion diversity.
Figure 5: Qualitative comparison of the state-of-the-art methods. Quantitative evaluations demonstrate that BioMoDiffuse achieves superior language-motion alignment, enhanced temporal smoothness, and improved physical plausibility compared to current approaches.
4.3 Ablation Study
Table 3: Ablation Study on traditional metrics. we evaluate the impact of adding $L_{accel}$ loss, incorporating the Physical Network, and introducing Physics Guidance on the performance of our method. All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables.
@@ 4.3 Ablation Study
$\checkmark$ $\times$ $\times$ $0.129$ $2.835$ $\bm{9.659}$ $1.882$ $0.537$ $0.734$ $0.827$
$\checkmark$ $\checkmark$ $\times$ $\bm{0.066}$ $\bm{2.783}$ $9.596$ $\bm{1.948}$ $\bm{0.549}$ $\bm{0.746}$ $\bm{0.836}$
@@
Table 4: Ablation Study on physical plausible metrics.
@@
$\checkmark$ $\times$ $\times$ $0.144$ $1.022$ $0.437$ $1.24$ $0.228$ $0.076$ $0.052$ $0.003$
$\checkmark$ $\checkmark$ $\times$ $0.135$ $\bm{0.954}$ $0.421$ $1.39$ $0.227$ $0.060$ $0.026$ $0.002$
$\checkmark$ $\checkmark$ $\checkmark$ $\bm{0.135}$ $0.959$ $\bm{0.385}$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $\bm{0.002}$
Our method introduces three key innovations to the diffusion framework: acceleration loss, a physics network grounded in Euler-Lagrange equations, and physics-guided sampling. Ablation studies under identical training settings Tab. 3 and Tab. 4 validate their contributions. While the acceleration loss enhances kinematic feature extraction by expanding the motion representation dimension from 263 to 329, it initially degrades generation quality (e.g., FID deteriorates by 43.3%) due to increased solution space complexity. The proposed physics network addresses this by embedding biomechanical constraints, reformulating motion generation as a constrained optimization problem. This reduces spatial complexity while enforcing rigid-body dynamics, leading to a 48.8% FID improvement with only a 6.68% parameter increase over the UNet baseline. Crucially, the network’s Euler-Lagrange foundation enables simultaneous gains in physical plausibility and generation quality. Finally, physics-guided sampling prioritizes biomechanical fidelity during denoising, stabilizing acceleration fitting and further improving physical metrics, albeit with a slight trade-off in generation scores—a phenomenon consistent with CFG principles.
4.4 Motion Refining Results
Table 5: Comparative results for pose refinement. We trained and tested the WHAM [60] baseline and WHAM $+$ our method using the official code on the 3DPW [70] and RICH [21] datasets. Bold numbers denote the most accurate method in each column. Accel is in $m/s^{2}$, all other errors are in $mm$.
@@ 4.4 Motion Refining Results
3DPW WHAM $36.34$ $61.46$ $70.64$ $6.52$
+ Inversion $\bm{36.13}$ $\bm{61.11}$ $\bm{70.52}$ $\bm{5.55}$
RICH WHAM $50.67$ $83.00$ $95.32$ $5.35$
+ Inversion $\bm{50.60}$ $\bm{81.80}$ $\bm{94.84}$ $\bm{4.06}$
As shown in Tab. 5, we validate our model’s motion refinement capability using WHAM [60], a state-of-the-art pose estimation framework that leverages temporal modeling to achieve leading results on metrics like acceleration error. By integrating our physics-aware network with WHAM via diffusion inversion, we significantly improve motion smoothness while also advancing pose prediction accuracy to new state-of-the-art levels. This indicates that our model, equipped with diffusion inversion, can be applied after any pose estimation module to refine the motion sequence, resulting in more fluent and physically plausible results. These results demonstrate that our physical network effectively enhances temporal continuity and overall robustness, even without acceleration loss supervision. This dual capability underscores the versatility of our approach, showcasing its ability to refine both generated and estimated motions by enforcing physical plausibility at the intersection of kinematics and dynamics.
4.5 Ablation Experiments on Speed Control Signal
Table 6: Ablation Study on weight of Speed Control Signal. We evaluate the impact of different Speed Control Signal on generation quality and physical plausibility metrics. All experiments use the HumanML3D dataset.
@@ 4.5 Ablation Experiments on Speed Control Signal
Our physical network incorporates a speed control signal within the autoregressive block. Experiments reveal that applying this signal only during inference (via input weighting) dynamically scales motion amplitude, with stronger signals producing larger movements. However, three critical observations emerge: (1) Traditional generation metrics (e.g., FID) remain unaffected by signal strength, aligning with the convention that model parameters require consistency between training and testing; (2) Certain physical metrics, such as ground penetration, foot clipping, and velocity/acceleration consistency, also show minimal sensitivity; (3) Metrics tied to acceleration magnitude—smoothness and skating ratio—decrease with stronger signals, while ground floating increases. We attribute this to the explicit acceleration modeling in our framework: since the network learns absolute acceleration values (rather than relative relationships), scaling the input signal alters acceleration magnitudes, thereby directly impacting physics-based metrics. This demonstrates our model’s unique ability to modulate motion amplitude through acceleration control while preserving generation quality—a trade-off avoided by conventional methods. Furthermore, these findings guide targeted optimization of physical plausibility by tuning acceleration-related parameters.
5 Conclusion
In this paper, we present BioMoDiffuse, a motion generation model built upon a physical network derived from extended Euler-Lagrange equations. This model incorporates the physical network into pose estimation using DDIM Inversion. BioMoDiffuse combines physics guidance with a state-of-the-art UNet backbone architecture, while simultaneously learning from electromyographic (EMG) signals and acceleration data. Experimental results indicate that BioMoDiffuse surpasses previous models in both generation quality and physical plausibility. However, due to current dataset limitations, BioMoDiffuse has not yet been tested on higher-frame-rate datasets with more comprehensive annotations. We leave it for future exploration.
Supplementary Material
Video:
We have provided supplemental videos in our supplemental material. In the supplementary video, we demonstrate the following: 1) the differences between the human skeletons generated by our method and those by StableMoFusion [22], 2) the differences in the skinned human models generated by our method and StableMoFusion, and 3) the text prompts used to generate these motions. All generations were produced using the same code under identical random seed conditions (seed=0). We reconstructed the meshes following the rendering approach provided by MLD [5].
Code:
The code has also been included in the supplementary materials, and detailed execution instructions can be found in the accompanying README file.
Appendix A Evaluation Metrics
In the design of physical plausibility metrics, we refer to the design of the Skating Ratio in Gmd [29], as well as the designs for Floating, Ground Penetration, and Foot Clipping in ReinDiffuse [18], as well as the Speed and Smoothness metrics provided by Fischer [10]. The specific definitions of these metrics are as follows:
  • Skating Ratio: Foot skating refers to the situation where a person’s foot is in contact with the ground but still has velocity. Given a human motion $\Theta$, the Foot Skating Ratio for the $i$-th frame can be expressed as:
    $FS^{i}(\Theta)=\exp\left(-||\dot{\Theta^{i}_{ft}}\cdot f^{i}_{S}\cdot f^{i}_{V}\cdot f^{i}_{\bar{V}}||_{2}\right)$ (14)
    where $\dot{\Theta^{i}_{ft}}$, $f^{i}_{S}$, $f^{i}_{V}$, and $f^{i}_{\bar{V}}$ represent the foot velocity, the contact label, the instantaneous velocity threshold of the foot, and the average velocity threshold of the foot at the $i$-th frame, respectively.
    $Float^{i}(\Theta)=\exp\left(-||(\Theta^{i}_{h}-h_{ground})\cdot f^{i}_{F}||_{2}\right)$ (15)
    where $\Theta^{i}_{h}$ denotes the y-axis coordinate of the lowest point of the action at the $i$-th frame, $h_{ground}$ is a threshold hyperparameter, set to 5cm in practice, and $f^{i}_{F}$ is a flag that equals 1 when $\Theta^{i}_{h}>h_{ground}$ and 0 otherwise.
    $GP^{i}(\Theta)=\exp\left(-||(h_{ground}-\Theta^{i}_{h})\cdot f^{i}_{P}||_{2}\right)$ (16)
    where $f^{i}_{P}$ is a flag that equals 1 when $\Theta^{i}_{h}<h_{ground}$ and 0 otherwise.
    $FC^{i}(\Theta)=\exp\left(-||(\Theta^{i}_{lf}-\Theta^{i}_{rf})\cdot f^{i}_{C}||_{2}\right)$ (17)
    where $\Theta^{i}_{lf}$ and $\Theta^{i}_{rf}$ are the joint coordinates of the left and right feet, and $f^{i}_{C}$ is a threshold function. $f^{i}_{C}=1$ when the distance between the two feet exceeds a threshold, which is set to 5cm in practice.
    $Speed^{i}(\Theta)=\frac{||\Theta^{i}||_{2}}{\Delta_{t}}$ (18)
  • Smoothness: Smoothness can be used to measure the jittering of the model’s acceleration. Models with poor quality tend to exhibit poor Smoothness. The calculation for Smoothness at the $i$-th frame is given by:
    $Smooth^{i}(\Theta)=\frac{|\ddot{\Theta^{i}}-\ddot{\Theta^{i-1}}|}{\Delta_{t}}$ (19)
It is important to note that all of the aforementioned metrics should ideally be as close as possible to the dataset. Since the dataset includes motions such as crawling, climbing stairs, and freestyle swimming, metrics like Skating Ratio and Floating are not necessarily better when smaller. For the KIT-ML dataset [51], Ground Penetration and Clipping are also not zero, which may be related to the quality of the KIT-ML dataset.
Appendix B Implementation Details

### v2

@@ III-D Generation with Diffusion Inversion
Figure 3: Estimation, refinement, and editing using DDIM Inversion. We incorporate the BioVAE-Diffusion framework as a residual module and apply diffusion inversion to enable pose estimation, motion refinement, and motion editing tasks. By adopting an inversion-then-forward pipeline, our approach significantly improves the smoothness of generated SMPL sequences while enhancing the overall generation quality.
@@
BioVAE can be readily adapted to pose estimation, motion refinement, and motion editing tasks. The key insight is that our BioVAE-Diffusion framework is capable of generating physically plausible motions starting from latent noise. Given the initial estimates of human motion, we first employ diffusion inversion [44, 46]—a technique that reverses the usual diffusion process—to recover the latent noise corresponding to the motion sequence. We then feed this noise into our model, which injects biomechanical constraints to produce motion sequences that are more authentic compared to the original inputs.
Specifically, given the initial SMPL [68] parameters $\mathbf{x}_{0}$, we first invert them to timestep $t$ using DDIM sampling:
@@
where $t$ controls the balance between information preservation and noise introduction. The inverted latent $\mathbf{x}_{t}$ then undergoes forward BioVAE-Diffusion to produce the refined motion:
@@
Crucially, our parametric implementation inherits the physics constraints from Sections III-B and III-C without requiring additional tuning, demonstrating the strong generalization capability of our method across multiple tasks.
TABLE II: Quantitative evaluation of physical plausible metrics on the HumanML3D and KIT-ML test set. All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables. The results for T2M, and MotionDiffuse on HumanML3D is sourced from ReinDiffuse [22].
@@
MoMask $0.106$ $1.048$ $0.331$ $\bm{1.89}$ $0.261$ $0.050$ $0.013$ $0.001$
Ours $\bm{0.135}$ $\bm{0.959}$ $0.385$ $1.45$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $0.002$
IV Experiments
IV-A Datasets and Metrics
We trained and evaluated our model on both the KIT-ML [16] and HumanML3D [15] datasets. The KIT-ML dataset contains 3,911 motions and 6,363 natural language annotations, while HumanML3D is derived from and reannotates the HumanAct12 [69] and AMASS [70] datasets, comprising a total of 14,616 motion sequences and 44,970 descriptions.
We follow the evaluation metrics used in previous motion generation works [24, 61, 64], including Frechet Inception Distance (FID), Precision, Diversity, Multimodality, and Multi-Modal Distance. Additionally, we consider physics-related metrics such as skating ratio, ground floating, ground penetration, and foot clipping, which have also been used in prior studies [22, 25, 23]. Furthermore, we introduce two novel metrics derived from Fourier analysis: Velocity Consistency (VC) and Acceleration Consistency (AC). Our motivation stems from the observation that while the Smoothness metric captures global temporal coherence, it overlooks the inherent kinematic relationship between stable body parts and agile extremities in human motion. For a motion sequence represented as $\Phi\in\mathbb{R}^{T\times 3d}$, the metrics are formally defined as:
IV-B Evaluation Metrics
In designing physical plausibility metrics, we draw on the Skating Ratio proposed in Gmd [26], the Floating, Ground Penetration, and Foot Clipping metrics from ReinDiffuse [22], and the Speed and Smoothness metrics introduced by Fischer [14]. The specific definitions of these metrics are as follows:
  • Skating Ratio: Foot skating refers to the situation where a person’s foot is in contact with the ground but still has velocity. Given a human motion $X$, the Foot Skating Ratio for the $i$-th frame can be expressed as:
    $FS^{i}(X)=\exp\left(-||\dot{X^{i}_{ft}}\cdot f^{i}_{S}\cdot f^{i}_{V}\cdot f^{i}_{\bar{V}}||_{2}\right),$ (13)
    where $\dot{X^{i}_{ft}}$, $f^{i}_{S}$, $f^{i}_{V}$, and $f^{i}_{\bar{V}}$ represent the foot velocity, the contact label, the instantaneous velocity threshold of the foot, and the average velocity threshold of the foot at the $i$-th frame, respectively.
    $Float^{i}(X)=\exp\left(-||(X^{i}_{h}-h_{ground})\cdot f^{i}_{F}||_{2}\right),$ (14)
    where $X^{i}_{h}$ denotes the y-axis coordinate of the lowest point of the action at the $i$-th frame, $h_{ground}$ is a threshold hyperparameter, set to 5cm in practice, and $f^{i}_{F}$ is a flag that equals 1 when $X^{i}_{h}>h_{ground}$ and 0 otherwise.
    $GP^{i}(X)=\exp\left(-||(h_{ground}-X^{i}_{h})\cdot f^{i}_{P}||_{2}\right),$ (15)
    where $f^{i}_{P}$ is a flag that equals 1 when $X^{i}_{h}<h_{ground}$ and 0 otherwise.
    $FC^{i}(X)=\exp\left(-||(X^{i}_{lf}-X^{i}_{rf})\cdot f^{i}_{C}||_{2}\right),$ (16)
    where $X^{i}_{lf}$ and $X^{i}_{rf}$ are the joint coordinates of the left and right feet, and $f^{i}_{C}$ is a threshold function. $f^{i}_{C}=1$ when the distance between the two feet exceeds a threshold, which is set to 5cm in practice.
    $Speed^{i}(X)=\frac{||X^{i}||_{2}}{\Delta_{t}},$ (17)
  • Smoothness: Smoothness serves as a measure of fluctuations in motion acceleration. Models of inferior quality often exhibit low smoothness in their generated motions. The calculation for Smoothness at the $i$-th frame is given by:
    $Smooth^{i}(X)=\frac{|\ddot{X^{i}}-\ddot{X^{i-1}}|}{\Delta_{t}}.$ (18)
It is important to note that all the aforementioned metrics should ideally be as close as possible to the data set. Since the text-to-motion datasets include motions such as crawling, climbing stairs, and swimming, metrics like Skating Ratio and Floating are not necessarily better when smaller. For the KIT-ML dataset [16], Ground Penetration and Clipping are also not zero, which may be related to the quality of the KIT-ML dataset.
IV-C Quantitative Results
Figure 4: User Study Results.
User Study.
We conducted a user study in which participants rated four aspects of the generated motions on a scale of 1 to 5: “Generation Quality”, “Motion Smoothness”, “Physical Plausibility” and “Text Alignment”, using the HumanML3D test set. Thirty participants were invited to evaluate and compare the motion generation results of StableMoFusion + BioVAE, StableMoFusion, and the ground truth across 20 prompts. The results show that our model outperforms StableMoFusion in generation quality, motion smoothness, physical plausibility, and semantic alignment. In particular, our method is capable of generating results nearly indistinguishable from the ground truth in some cases, highlighting its effectiveness.
Figure 5: User study interface for evaluating on HumanML3D. To ensure unbiased feedback, the samples were presented in random order to avoid potential scoring inertia among participants.
@@ IV-C Quantitative Results
Table I and Table II compare our framework with existing approaches in both physical plausibility metrics (e.g., smoothness, foot sliding) and traditional generation metrics (FID, Diversity). Our approach achieves state-of-the-art results in biomechanical authenticity while simultaneously advancing traditional motion quality, outperforming prior works across multiple dimensions. Experiments reveal critical limitations in existing paradigms: GPT-style generative models tend to over-smooth motions, failing to capture realistic dynamics, while diffusion-based methods exhibit temporal discontinuity, probably due to their implicit treatment of spatio-temporal relationships without explicit temporal modeling. In contrast, our method leverages the residual joint training of BioVAE and diffusion models, enabling the diffusion process to effectively capture acceleration information during training. Meanwhile, the incorporation of biomechanical priors reduces the complexity of the solution space, allowing our model to achieve superior performance on traditional metrics without compromising motion diversity.
Figure 6: Qualitative comparison of the state-of-the-art methods. The results demonstrate that our method achieves superior language-motion alignment, enhanced temporal smoothness, and improved physical plausibility compared to current approaches.
IV-D Ablation Study
TABLE III: Ablation study on traditional metrics. All experiments were repeated 20 times with different random seeds, and the mean results are reported in the tables.
@@ IV-D Ablation Study
$\checkmark$ $\times$ $\times$ $0.085$ $\bm{2.760}$ $9.764$ $1.835$ $\bm{0.551}$ $\bm{0.747}$ $\bm{0.839}$
$\checkmark$ $\checkmark$ $\times$ $\bm{0.066}$ $2.783$ $9.596$ $\bm{1.948}$ $0.549$ $0.746$ $0.836$
$\checkmark$ $\times$ $\checkmark$ $0.077$ $2.803$ $9.752$ $1.873$ $0.545$ $0.740$ $0.832$
@@
TABLE IV: Ablation study on physical plausible metrics.
@@
$\checkmark$ $\times$ $\times$ $0.161$ $1.205$ $0.418$ $1.23$ $0.227$ $0.056$ $0.026$ $0.002$
$\checkmark$ $\checkmark$ $\times$ $0.135$ $0.954$ $0.421$ $1.39$ $0.227$ $0.060$ $0.026$ $0.002$
$\checkmark$ $\times$ $\checkmark$ $\bm{0.134}$ $\bm{0.944}$ $0.425$ $1.33$ $0.211$ $0.056$ $0.039$ $0.002$
$\checkmark$ $\checkmark$ $\checkmark$ $0.135$ $0.959$ $0.385$ $\bm{1.45}$ $\bm{0.204}$ $\bm{0.056}$ $\bm{0.013}$ $\bm{0.002}$
Our proposed BioVAE not only incorporates additional biomechanical priors but also achieves stable joint training by modifying the diffusion loss function. Ablation studies in identical training settings (Table III and Table IV) demonstrate the critical contributions of the BioVAE architecture, the biomechanical priors, and the newly introduced loss term. Specifically, BioVAE substantially improves motion generation performance on traditional metrics, while the biomechanical priors and BioVAE loss each independently enhance the physical plausibility of the generated motions. When combined, these components exhibit even greater improvements, highlighting the effectiveness and advancement of our framework.
Figure 7: Additional samples from our text-to-motion synthesis approach, generated with text prompts from the HumanML3D test set. The results demonstrate that our approach effectively understands diverse textual semantics and generates physically plausible motions.
IV-E Motion Estimation and Editing
TABLE V: Comparative results for pose estimation. We used WHAM [4] and various diffusion models as the base framework. Building on the official WHAM implementation, we evaluated the impact of incorporating the residual-based BioVAE on pose estimation performance using the 3DPW [17] and RICH [18] datasets. Bold numbers denote the most accurate method in each column. Accel is in $m/s^{2}$, all other errors are in $mm$.
Method 3DPW RICH
@@ IV-E Motion Estimation and Editing
WHAM 36.34 61.46 70.64 6.52 50.67 83.00 95.32 5.35
MDM Inversion 36.48 62.06 71.36 6.41 51.76 83.62 96.53 5.28
StableMoFusion Inversion $\bm{36.06}$ 61.85 71.26 6.29 51.64 83.12 96.53 5.11
StableMoFusion + BioVAE Inversion 36.13 $\bm{61.11}$ $\bm{70.52}$ $\bm{5.55}$ $\bm{50.60}$ $\bm{81.80}$ $\bm{94.84}$ $\bm{4.06}$
As shown in Table V, we evaluate the capability of our framework for motion refinement and pose estimation using WHAM [4], a state-of-the-art method in the pose estimation domain. WHAM incorporates sequential modeling and temporal information, achieving leading results in acceleration-related metrics. Our results demonstrate that integrating BioVAE with diffusion models and pose estimation frameworks can significantly improve the smoothness of human motion while also enhancing pose estimation accuracy, achieving state-of-the-art performance across all evaluated metrics.
These experiments indicate that our physics-aware network not only reduces the complexity of the solution space and enhances the physical plausibility but also substantially improves the smoothness of the motion and the overall performance even without relying on acceleration loss supervision. This highlights the versatility of our approach to simultaneously refining human motion and advancing pose estimation results.
TABLE VI: Motion Editing via Acceleration Modification. We manually edited the acceleration values in BioVAE to evaluate the effects of increasing or decreasing acceleration on the generated motions. All experiments were conducted on the HumanML3D dataset.
@@
By modifying the acceleration values in BioVAE for ablation studies, we observed the following findings: (1) Traditional generation metrics generally do not improve with changes in acceleration magnitude, which explains why generative models require parameter consistency between training and testing; (2) Certain physical metrics, such as ground penetration, foot clipping, and velocity/acceleration consistency, also show minimal sensitivity to acceleration changes; (3) Some physical metrics exhibit a strong correlation with acceleration magnitude: smoothness and skating ratio decrease as acceleration increases, while ground floating increases with higher acceleration values.
We attribute these phenomena to the explicit representation of acceleration in our model. When we modify the absolute magnitude of the input acceleration, the model shows little impact on traditional generation metrics. However, because our method explicitly models acceleration, altering its absolute magnitude affects the generated motion amplitude and, in turn, influences acceleration-related physical plausibility metrics. This highlights the advantage of our approach: the ability to regulate acceleration to control the physical properties of generated motions. In addition, these insights provide guidance for optimizing specific physical metrics by tuning acceleration-related parameters.
IV-F Implementation Details